eviant:

Solve the logarithmic equation by writing in exponent form: log 5x + 3 = 7

eviant:

@dude @Nnesha

justjm:

\(\log5x+3=7\) \(\log5x=4\) Use your understanding that: \(a^{b}=c\) \(\log_{a}c=b\) are the same. I am using this approach because it says "...by writing in exponent form" \(10^{4}=5x\) Now you can easily solve for x.

jhonyy9:

may be (2*5)^4 =5x (2^4)*(5^4) = 5x divide both sides by 5 so what you get ?